# Hessian estimates for Lagrangian mean curvature equation with Lipschitz critical and supercritical phases

### v1

@@ 1. Introduction
It’s recorded as (2.1) in [61], solutions to (1.1) for supercritical phases are semi-convex, and (in principle) have better ’compactness’, compared with solutions to (1.1) for critical and supercritical phases. Another important advantage of the supercritical condition on the phase $\theta$ is that it can give a Jacobi inequality involving a divergence term of derivatives of $\theta$ (see [64]), while such a Jacobi inequality does not hold for critical and supercritical phases.
@@
We will use the symbol $\ell$ to represent the subscript in sequences of functions $u_{\ell}$ or $\theta_{\ell}$, which does not mean the derivative of $u$ or $\theta$ w.r.t. $x_{\ell}$. We use $u_{i},w_{i}$ to denote the derivatives of $u,w$ w.r.t. $x_{i}$ for English letter $i$ (and similar ones for $j,k,l$). For an open $\Omega\subset\mathbb{R}^{n}$ and a function $u\in C^{2}(\Omega)$, we always use $G_{Du}$ to denote a (Lagrangian) graph of $Du$ over $\Omega$ in $\mathbb{R}^{n}\times\mathbb{R}^{n}$. For a function $f$ on $G_{Du}$, we will identify $f(x,Du(x))=f(x)$, and regard it as a function $f$ on $\Omega$ in context, and vice versa.
@@ 2.1. Geometry of Lagrangian graphs in Euclidean space
where the equality occurs if and only if $\mathbf{v}$ represents a Lagrangian plane with the phase $\theta_{L}(\mathbf{p})$ mod $2\pi$. Hence, $\theta=\theta_{L}$ mod $2\pi$. For the constant $\theta$, $L$ is minimal, and $L$ is said to be a minimal (or special) Lagrangian graph (over $\Omega$). This is also equivalent to that $L$ (or equivalently the current associated with $L$) is minimizing in $\Omega\times\mathbb{R}^{n}$ (see Theorem 2.3, Proposition 2.17 in [31]; or [59]).
@@
(Noting that $\lambda_{i}$ may be not $C^{1}$, and $\gamma$ may be not $C^{0}$.) Let $\theta_{\gamma_{i}}=D_{\gamma_{i}}\theta=\langle\gamma_{i},D\theta\rangle$ and
@@
for each $i=1,\cdots,n$. (Noting $e_{i},\nu_{i}$ may be not $C^{0}$.) At any considered point $\mathbf{p}=(p,Du(p))$, by using tangent vector fields $\gamma_{i}(p)+\sum_{j=1}^{n}\langle\gamma_{i}(p),(D^{2}u)\gamma_{j}(p)\rangle J\gamma_{j}(p)$ on $L$, we get
@@ 2.2. Varifolds and currents in geometric measure theory
then there is a function $f=(f_{1},\cdots,f_{N-n})$ on a ball $\mathbf{B}_{\varrho_{*}r}\cap P$ for some $n$-subspace $P\subset\mathbb{R}^{N}$ so that $\{(x,f(x)):\,x\in B_{\varrho_{*}r}\}\subset\mathrm{spt}V$ and
@@ 3. Integral estimates for Lagrangian graphs
In this section, we will establish two integral estimates for Lagrangian graphs mainly based on Wang-Yuan’s results in [54]. Let $n\geq 2$, and $(u,\theta)\in\mathbb{F}_{n}(\Lambda,\kappa,r)$ for some $\Lambda\geq 0$, $\kappa\geq 1$ and $r>0$. Let $\delta=\inf_{B_{r/2}}\theta-\frac{n-2}{2}\pi$. For $\delta>0$, let us consider Lewy-Yuan rotation in [60] as follows. Let $\bar{x},\bar{y}:\ B_{r}\to\mathbb{R}^{n}$ be smooth mappings defined by
@@
If $\delta\geq\pi/2$, then $u$ is convex, and $-\tan\frac{\pi}{2n}\leq D^{2}\bar{u}\leq\cot\frac{\pi}{2n}$ from (9.10), which also gives (3.22). By considering a suitable finite covering $\{B_{1/2}(p_{i})\}$ of $B_{6}$, we complete the proof. ∎
@@ 4. Allard’s regularity theorem and Harnack’s inequality
From (4.14), we get that $\hat{\lambda}_{1}$ has multiplicity one w.r.t. $D^{2}w$ on $B_{\delta}$. From Proposition 4.1 and $\gamma_{n}(0^{n})=E_{n}^{T}$, we can choose the suitably small $\delta>0$ so that $|\gamma_{n}-E_{n}|$ is sufficiently small on $B_{\delta}$, which implies
@@ 5. Hessian estimates with constraints
$=$ $\lambda_{k}^{2}\left(\sum_{i=1}^{n-1}h_{iik}\right)^{2}\left(\sum_{i\leq n-1}\frac{\lambda_{k}}{2\lambda_{i}}+\frac{1}{3}\right)^{-1}+\frac{\lambda_{k}^{2}}{\frac{\lambda_{k}}{2\lambda_{n}}-\frac{1}{3}}h_{nnk}^{2}.$
@@
$=$ $\sum_{l=m+1}^{n}\sum_{i=1}^{n}\frac{h_{iil}\theta_{\gamma_{l}}}{\sqrt{1+\lambda_{l}^{2}}}-\sum_{l=m+1}^{n}\sum_{i\neq l}^{n}\frac{\lambda_{i}}{\lambda_{i}-\lambda_{l}}\sqrt{1+\lambda_{l}^{2}}h_{iil}\theta_{\gamma_{l}}$
$+\sum_{i=m+1}^{n}\sum_{l\neq i}^{n}\frac{\lambda_{i}}{\lambda_{i}-\lambda_{l}}\sqrt{1+\lambda_{l}^{2}}h_{iil}\theta_{\gamma_{l}}$
@@ 6. A rigidity theorem for special Lagrangian submanifolds
$\cos\Theta_{i}(z)\leq\Theta_{i}(z_{0})-\Theta_{i}(z)\leq\psi_{\Theta}(|z-z_{0}|)|z-z_{0}|\qquad\mathrm{for\ each}\ i\in\{1,\cdots,m\}.$
So we have
@@ 8. Application to Dirichlet problem of Lagrangian mean curvature equation
There are infinite sequences of smooth functions $\{\theta_{\ell}\}_{\ell\geq 1}$ and $\{\psi_{\ell}\}_{\ell\geq 1}$ so that $\theta_{\ell}:\overline{\Omega}\,\to[\frac{n-2}{2}\pi+\delta_{\ell},\frac{n}{2}\pi)$ for a constant $\delta_{\ell}>0$ with $\lim_{\ell\to\infty}\delta_{\ell}=0$, $\lim_{\ell\to\infty}\theta_{\ell}=\theta$, $\lim_{\ell\to\infty}D(\theta_{\ell}-\theta)=0$ a.e., and $\psi_{\ell}\to\psi$ in $C^{1,1}$-sense on $\partial\Omega$. From Theorem 4.1 in [6], there exists a unique solution $u_{\ell}\in C^{2}(\overline{\Omega})$ to

### v2

@@ 1. Introduction
It’s recorded as (2.1) in [61], solutions to (1.1) for supercritical phases are semi-convex, and (in principle) have better ’compactness’, compared with solutions to (1.1) for critical and supercritical phases. Another important advantage of the supercritical phase $\theta$ is that it can give a Jacobi inequality involving a divergence term of derivatives of $\theta$ (see [64]), while such a Jacobi inequality does not hold for critical and supercritical phases.
@@
We will use the symbol $\ell$ to represent the subscript in sequences of functions $u_{\ell}$ or $\theta_{\ell}$, rather than the derivative of $u$ or $\theta$ w.r.t. $x_{\ell}$. We use $u_{i},w_{i}$ to denote the derivatives of $u,w$ w.r.t. $x_{i}$ for English letter $i$ (and similar ones for $j,k,l$). For an open $\Omega\subset\mathbb{R}^{n}$ and a function $u\in C^{2}(\Omega)$, we always use $G_{Du}$ to denote a (Lagrangian) graph of $Du$ over $\Omega$ in $\mathbb{R}^{n}\times\mathbb{R}^{n}$. For a function $f$ on $G_{Du}$, we will identify $f(x,Du(x))=f(x)$, and regard it as a function $f$ on $\Omega$ in context, and vice versa.
@@ 2.1. Geometry of Lagrangian graphs in Euclidean space
where the equality occurs if and only if $\mathbf{v}$ represents a Lagrangian plane with the phase $\theta_{L}(\mathbf{p})$ mod $2\pi$. Hence, $\theta=\theta_{L}$ mod $2\pi$. For the constant $\theta$, $L$ is minimal, and $L$ is said to be a minimal (or special) Lagrangian graph (over $\Omega$). This is also equivalent to that (the current associated with) $L$ is minimizing in $\Omega\times\mathbb{R}^{n}$ (see Theorem 2.3, Proposition 2.17 in [31]; or [59]).
@@
Noting that $\lambda_{i}$ may not be $C^{1}$, and $\gamma$ may not be $C^{0}$. Let $\theta_{\gamma_{i}}=D_{\gamma_{i}}\theta=\langle\gamma_{i},D\theta\rangle$ and
@@
for each $i=1,\cdots,n$. At any considered point $\mathbf{p}=(p,Du(p))$, by using tangent vector fields $\gamma_{i}(p)+\sum_{j=1}^{n}\langle\gamma_{i}(p),(D^{2}u)\gamma_{j}(p)\rangle J\gamma_{j}(p)$ on $L$, we get
@@ 2.2. Varifolds and currents in geometric measure theory
then there is a function $f=(f_{1},\cdots,f_{N-n})$ on a ball $\mathbf{B}_{\varrho_{*}r}\cap P$ for some $n$-subspace $P\subset\mathbb{R}^{N}$ so that $\{(x,f(x)):\,x\in\mathbf{B}_{\varrho_{*}r}\cap P\}\subset\mathrm{spt}V$ and
@@ 3. Integral estimates for Lagrangian graphs
In this section, we will establish two integral estimates for Lagrangian graphs mainly based on Wang-Yuan’s results in [54]. Let $n\geq 2$, and $(u,\theta)\in\mathbb{F}_{n}(\Lambda,\kappa,r)$ for some $\Lambda\geq 0$, $\kappa\geq 1$ and $r>0$. Let $\delta=\inf_{B_{r/2}}\theta-\frac{n-2}{2}\pi$. For $\delta>0$, let us consider Lewy-Yuan rotation in [60] as follows. Let $\bar{x},\bar{y}:\ B_{r}\to\mathbb{R}^{n}$ be smooth mapping defined by
@@
If $\delta\geq\pi/2$, then $u$ is convex, and $-\tan\frac{\pi}{2n}\leq D^{2}\bar{u}\leq\cot\frac{\pi}{2n}$ from (9.10) in Appendix I, which also gives (3.22). By considering a suitable finite covering $\{B_{1/2}(p_{i})\}$ of $B_{6}$, we complete the proof. ∎
@@ 4. Allard’s regularity theorem and Harnack’s inequality
From (4.14), we get that $\hat{\lambda}_{1}$ has multiplicity one w.r.t. $D^{2}w$ on $B_{\delta}$. From Proposition 4.1 and $\gamma_{n}(0^{n})=E_{n}^{T}$, we can choose the suitably small $\delta>0$ so that $|\gamma_{n}-E_{n}|$ is sufficiently small on $B_{\delta}$. Then it can give
@@ 5. Hessian estimates with constraints
$\geq$ $\lambda_{k}^{2}\left(\sum_{i=1}^{n-1}h_{iik}\right)^{2}\left(\sum_{i\leq n-1}\frac{\lambda_{k}}{2\lambda_{i}}+\frac{1}{3}\right)^{-1}+\frac{\lambda_{k}^{2}}{\frac{\lambda_{k}}{2\lambda_{n}}-\frac{1}{3}}h_{nnk}^{2}.$
@@
$=$ $\sum_{l=m+1}^{n}\sum_{i=1}^{n}\frac{h_{iil}\theta_{\gamma_{l}}}{\sqrt{1+\lambda_{l}^{2}}}-\sum_{l=m+1}^{n}\sum_{i\neq l}\frac{\lambda_{i}}{\lambda_{i}-\lambda_{l}}\sqrt{1+\lambda_{l}^{2}}h_{iil}\theta_{\gamma_{l}}$
$+\sum_{i=m+1}^{n}\sum_{l\neq i}\frac{\lambda_{i}}{\lambda_{i}-\lambda_{l}}\sqrt{1+\lambda_{l}^{2}}h_{iil}\theta_{\gamma_{l}}$
@@ 6. A rigidity theorem for special Lagrangian submanifolds
$\cos\Theta_{i}(z)=\cos\Theta_{i}(z)-\cos\frac{\pi}{2}\leq\frac{\pi}{2}-\Theta_{i}(z)=\Theta_{i}(z_{0})-\Theta_{i}(z)\leq\psi_{\Theta}(|z-z_{0}|)|z-z_{0}|$
for each $i\in\{1,\cdots,m\}$. So we have
@@ 8. Application to Dirichlet problem of Lagrangian mean curvature equation
There are two infinite sequences of smooth functions $\{\theta_{\ell}\}_{\ell\geq 1}$ and $\{\psi_{\ell}\}_{\ell\geq 1}$ so that $\theta_{\ell}:\overline{\Omega}\,\to[\frac{n-2}{2}\pi+\delta_{\ell},\frac{n}{2}\pi)$ for a constant $\delta_{\ell}>0$ with $\lim_{\ell\to\infty}\delta_{\ell}=0$, $\lim_{\ell\to\infty}\theta_{\ell}=\theta$, $\lim_{\ell\to\infty}D(\theta_{\ell}-\theta)=0$ a.e., and $\psi_{\ell}\to\psi$ in $C^{1,1}$-sense on $\partial\Omega$. From Theorem 4.1 in [6], there exists a unique solution $u_{\ell}\in C^{2}(\overline{\Omega})$ to